OpenStudy (anonymous):

(Law of Motion) A force of 300 n slows down a wagon from 40 kph to 20 kph in 10 seconds. What is the wagon's weight?

OpenStudy (anonymous):

Use the equation \[\Sigma F = ma\] and solve for m then after you have m you use \[F_g = mg\] where g = 9.8 N/kg

OpenStudy (anonymous):

Would that be the only formula for the problem?

OpenStudy (anonymous):

two different equations I am not replace Fg with the top equation

OpenStudy (anonymous):

okay thanks :)

OpenStudy (anonymous):

let me know if you need more assistance

OpenStudy (anonymous):

Could you help me in using ΣF=ma ? I dont know how to solve using this equation

OpenStudy (anonymous):

I believe the deceleration should be found first. using 1 of 3 kinematic equations, find what -acceleration is required to go from 40km/h down to 20km/h in 10 seconds. If your not sure which equation to use, let me know

OpenStudy (anonymous):

\[velocity= \frac{ \Delta displacement }{ time }\] is this correct?

OpenStudy (anonymous):

wait.. have you been introduced to these equations? v^2=v^2 + 2as v=v+at s=s+vt+(1/2)at^2

OpenStudy (anonymous):

no. :/

OpenStudy (anonymous):

they don't look familiar at all? none zip no-way no-how?

OpenStudy (anonymous):

they dont look familiar

OpenStudy (anonymous):

ok ^_^ no problem so we will find acceleration using this equation: acceleration = (velocity_2 - velocity_1)/time

OpenStudy (anonymous):

where velocity_2 is 40km/h velocity_ = 20km/h

OpenStudy (anonymous):

when you wrote, " 40 kph to 20 kph" do you mean "40km per hour to 20km per hour"?

OpenStudy (anonymous):

yeah :)

OpenStudy (anonymous):

ok good :) so anytime you do use an equation that has to do with velocity. we have to make sure that velociy is meters per second. here you are given kilometers per hour. so you'll need to convert km/h into m/s, can you do this?

OpenStudy (anonymous):

yeah. :) so the new one is 11.11 m/s to 5.56 m/s Am I right?

OpenStudy (anonymous):

yes, good work ^_^ now we will use those to figure out the negative acceleration that the wagon is experiencing with this: \[acceleration = \frac{ Velocity_{faster} - Velocity_{slower} }{ time}\]

OpenStudy (anonymous):

then the acceleration would be = \[\frac{ 20 m/s }{ 10s }\] ?

OpenStudy (anonymous):

then the result would be 2 m? is that an acceleration? or am I worng?

OpenStudy (anonymous):

think again.. remember what you got here: "yeah. :) so the new one is 11.11 m/s to 5.56 m/s Am I right?" use these numbers, becuase we have to have it in m/s

OpenStudy (anonymous):

ooops. i forgot so its \[\frac{ 5.55 m/s }{ 10 s}\]

OpenStudy (anonymous):

very good ^_^

OpenStudy (anonymous):

thanks :) so what's next?

OpenStudy (anonymous):

so finishing out the math, we have 5.55/10 = 0.55 0.55 m/s^2 is the rate at which the wagon is being slowed down (negative acceleration)

OpenStudy (anonymous):

now that we know how much it is being slowed down. we can find out it's mass. this is where we will use the equation: force = mass*acceration we know acceleration, we know force [given in the problem information] so now we solve for mass ^_^

OpenStudy (anonymous):

so, in finding mass, it would be \[\frac{ 300N }{ 0.55 m/s ^{2}}\] then the result would be 545.45 kg

OpenStudy (anonymous):

bingo ^_^

OpenStudy (anonymous):

Ohhhhhhhhhhhhhhh! Thanks so much :)

OpenStudy (anonymous):

your welcome ^_^ i'm logging off now, have a good evening

OpenStudy (anonymous):

You too. Thanks again :)